40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Aug 11 17:34:16 2017, awade, Craig, DailyProgress, BEAT, Mapping the beat notes north temp vs south temp NorthLaserTempVsPrecavityBeatnoteFrequency.pdf
    Reply  Sun Aug 13 19:58:49 2017, awade, Craig, DailyProgress, BEAT, Fixing the calibration of North laser Lightwave series 125/126 set point temperature 
       Reply  Mon Aug 14 16:34:45 2017, awade, Craig, DailyProgress, BEAT, Voltage offsets on unterminated slow frequency control on Lightwave lasers 2017-08-14_16.14.12.jpg2017-08-14_16.14.25.jpg
Message ID: 1857     Entry time: Sun Aug 13 19:58:49 2017     In reply to: 1855     Reply to this: 1858
Author: awade, Craig 
Type: DailyProgress 
Category: BEAT 
Subject: Fixing the calibration of North laser Lightwave series 125/126 set point temperature 

When Craig and I looked and mapping out good beat notes on Friday it wasn't clear that we had enough range to find more that one pair of temperatures that would work.  The two lasers seemed to have vastly different operating set points for a corresponding frequency match and re centering the slow controls BNC inputs to zero seemed to go in the opposite direction to that expect.  

As noted in the previous post summarizing config of the lasers (PSL:1854) South laser was set at +47.0749 C and North laser was set at +26.4650 C.  We were reluctant to tune the North laser much lower than 24 C, lest we risk condensation of water within the head. This meant that once the slow frequency controls voltage was reset back to zero, there was very little range of the North laser temperature settings left to find a BN within.

After checking in the manual, it turns out two things were not what we expected.  Firstly, the slow tuning has opposite sign to the laser temperature: increasing temperature tunes laser frequency down and increasing slow frequency voltage tunes laser frequency up. Secondly the laser set point temperature is a calculated value and note a true value (the screen value LT gives the true value).  It also doesn't include the offset caused by voltage on the front panel Slow frequency BNC.  So for North laser set point temperature of +26.4650 C the true value measured at the crystal is 44.8 C (with zero voltage on slow freq input).  

The discrepancy was probably caused by a poor calibration on setting up the laser head with the controller box.  The setup procedure when configuring a new laser head is to center the temperature setting in the middle of the range, wait for it to come to equilibrium and then to trigger a recalibration of temperature to this value.  This is done by putting the laser into standby and pressing: Display x2 (will read C +xx.xxxx) -> Set (to recenter real temperature) -> wait for equilibrium -> Set (to recenter the calculated set point temperature). 

It is likely that set was pressed twice in rapid succession and that the calibration was triggered before the laser had time to reach its midpoint temperature.  This hasn't affect operation, but it does mean that the set point looked much lower than it should have been, we might have been worried about selecting temperatures that were too low to safely operate the laser at. 

I have recalibrate the laser 'Center Temperature Calibration' on both lasers (North was too low and south was too high) so that we can have more confidence that the laser ranges are in a safe range.  Calibrated setpoints are now 48.4 ± 15 C for the South laser and 44.8 C ± 15 C for the North laser.

ELOG V3.1.3-